Vector A has magnitude 4.0 units and vector B has magnitude 6.0 units. The angle between A and B is 60.0°. What is the magnitude of 2A+3B ?



1
Expert's answer
2021-04-05T11:09:57-0400

Let's direct the x-axis of Cartesian coordinate along vector "\mathbf{A}". Then the coordinates of vector "\mathbf{A}" will be:


"\mathbf{A} = 4(1,0) = (4,0)"

The coondinates of vector "\mathbf{B}" are;


"\mathbf{B} = 6(\cos60\degree,\sin60\degree) = (3\sqrt{3},3)"

Then the combitanion will have the following coordinates:


"2\mathbf{A} + 3\mathbf{B} = 2(4,0) + 3(3\sqrt{3},3) = (8+9\sqrt3 , 9)"

The magnitude of this vector is:


"|2\mathbf{A} + 3\mathbf{B}| = \sqrt{(8+9\sqrt{3})^2 + 9^2} \approx 25.2"

Answer. 25.2.